# OpenMathInstruct-1: A 1.8 Million Math Instruction Tuning Dataset

### v1

@@ Abstract
Recent work has shown the immense potential of synthetically generated datasets for training large language models (LLMs), especially for acquiring targeted skills. Current large-scale math instruction tuning datasets such as MetaMathQA (Yu et al., 2024) and MAmmoTH (Yue et al., 2024) are constructed using outputs from closed-source LLMs with commercially restrictive licenses. A key reason limiting the use of open-source LLMs in these data generation pipelines has been the wide gap between the mathematical skills of the best closed-source LLMs, such as GPT-4, and the best open-source LLMs. Building on the recent progress in open-source LLMs, our proposed prompting novelty, and some brute-force scaling, we construct OpenMathInstruct-1, a math instruction tuning dataset with 1.8M problem-solution pairs. The dataset is constructed by synthesizing code-interpreter solutions for GSM8K and MATH, two popular math reasoning benchmarks, using the recently released and permissively licensed Mixtral model. Our best model, OpenMath-CodeLlama-70B, trained on a subset of OpenMathInstruct-1, achieves a score of 84.6% on GSM8K and 50.7% on MATH, which is competitive with the best gpt-distilled models. We release our code, models, and the OpenMathInstruct-1 dataset under a commercially permissive license. [^1] Data and models are available at https://huggingface.co/collections/nvidia/openmath-65c5619de2ba059be0775014
@@
spacing=nonfrench
OpenMathInstruct-1: A 1.8 Million Math Instruction Tuning Dataset
Shubham Toshniwal, Ivan Moshkov, Sean Narenthiran, Daria Gitman, Fei Jia, Igor Gitman NVIDIA
@@ 1 Introduction
The huge development and inference costs associated with general-purpose large language models (LLMs) have led to the rise of smaller, task-specific LLMs. Recent work has proposed creating these domain/task-specific LLMs by generating high-quality synthetic data using powerful closed-source models such as GPT-3.5/4 (OpenAI et al., 2023) and training smaller models on the generated distillation data (Eldan and Li, 2023; Gunasekar et al., 2023; Li et al., 2023). For mathematical reasoning, our task of interest, all the current state-of-the-art open-source models are gpt-distilled (Wang et al., 2024; Yue et al., 2024; Gou et al., 2024; Liao et al., 2024). However, model development recipes relying on proprietary models like GPT-4 can have serious limitations: (a) legal restraints on how the finetuned models can be used, [^2] https://openai.com/policies/terms-of-use (b) generating data with closed-source models is typically costlier than state-of-the-art open-source models, and (c) these recipes lack reproducibility as closed-source model behaviors can vary significantly over time (Chen et al., 2023a).
@@
Figure 1: Training set coverage of Mixtral model generated solutions as a function of number of solutions sampled per problem (using temperature of 1.0 and top_p = 0.95). The statistics for the training set coverage of GPT-4 are from Gou et al. (2024).
@@
For developing mathematical reasoning models, why are open-source models not used in place of closed-source models? To answer this, we compare GPT-4 with Mixtral 8x7B model (Jiang et al., 2024), currently one of the best open-source LLMs at mathematical reasoning, by generating code-interpreter style solutions for two popular mathematical reasoning benchmarks, namely GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). We use the metric training set coverage (TSC) to compare the models, where TSC measures the number of training problems for which any of the generated solutions leads to the ground truth answer (pass@k). Figure 1 shows the training set coverage (TSC) of the Mixtral model as a function of the number of sampled solutions. For the relatively easier GSM8K benchmark, the Mixtral model’s coverage catches up to GPT-4’s with almost 8x the number of solution samples. For the challenging MATH benchmark, even with 12x the number of solutions, the Mixtral model still has a lower TSC than GPT-4. This gap in the training set coverage reflects the distillation data quality and, hence, the quality of the final fine-tuned model. This explains the preference for GPT-4 in the current distillation pipelines for mathematical reasoning.
@@
Dataset Size Generating LM
(Permissive License)
Lila (Mishra et al., 2022) 272K - ()
MathInstruct (Yue et al., 2024) 262K GPT-4 ()
MetaMathQA (Yu et al., 2024) 395K GPT-3.5 ()
MathCodeInstruct (Wang et al., 2024) 180K GPT-4 + Self ()
WizardMath* (Luo et al., 2023) 196K GPT-3.5 ()
ToRA* (Gou et al., 2024) 116K GPT-4 ()
@@
Bridging the coverage gap between GPT-4 and Open-source LLMs: We limit our investigation of open-source LLMs for synthesizing solutions to the Mixtral-base model due to (a) its strong performance on mathematical reasoning tasks compared to other open-source LLMs and (b) its permissive license. [^3] https://mistral.ai/news/mixtral-of-experts/ As a first attempt, we use a brute-force approach of sampling several solutions per problem. However, this approach only scales logarithmically, limiting its effectiveness (Figure 1). Next, we explore the approach of targeted solution generation, where we write few-shot prompts focused on specific sections of the training data. Concretely, we write few-shot prompts for each mathematics subject in the MATH dataset and merge the synthesized solutions. The motivation is that these subject-specific few-shot prompts could better target the latent mathematical capabilities of these general-purpose LLMs. Unfortunately, we only find a marginal gain in TSC with this approach (Section 2.2.2). Finally, we utilize the fact that text solutions accompany mathematical benchmarks such as MATH and GSM8K. These text solutions can aid the synthesis of code-interpreter style solutions. We show that using the text solution in our few-shot prompt with a slight modification substantially increases the coverage and, consequently, the performance of the fine-tuned model (Section 2.2.3).
Our solution synthesis experiments result in OpenMathInstruct-1, a collection of 1.8M problem-solution pairs. OpenMathInstruct-1 has a training set coverage of 93% for MATH and 99.9% for GSM8K. Table 1 shows that compared to previous mathematical reasoning fine-tuning datasets, OpenMathInstruct-1 is at least four times bigger and, even more importantly, it is permissively licensed, allowing unrestricted usage by future work. To illustrate the quality of OpenMathInstruct-1, we train and release a range of models based on Mistral-7B (Jiang et al., 2023), Llama 2 Touvron et al. (2023), and CodeLlama (Rozière et al., 2023). In particular, the CodeLlama-70B model fine-tuned on a subset of OpenMathInstruct-1, referred to as OpenMath-CodeLlama-70B, achieves a score of 84.6% on GSM8K and 50.7% on MATH. These scores are competitive with the current best gpt-distilled models. Finally, to support the open-source efforts in this direction, we publicly release all our fine-tuned models, code, and the OpenMathInstruct-1 dataset along with a further 6.6M incorrect sampled solutions. [^4] The incorrect solution trajectories can be used to train verifier models (Cobbe et al., 2021; Yu et al., 2023; Lightman et al., 2023).
@@ 2.1 Overview
Let $\mathcal{X}=\{(q_{1},a_{1}),\cdots,(q_{N},a_{N})\}$ be a typical mathematical reasoning training dataset, where $q_{i}$ and $a_{i}$ denote the $i^{\text{th}}$ question and answer respectively. Optionally, the training data may include text solution $t_{i}$, which illustrates a trajectory from $q_{i}$ to $a_{i}$ using mathematical principles. [^5] Both GSM8K and MATH have these text solutions. Besides the data, we assume access to a foundation LLM like Mixtral-base. The goal is to generate diverse, high-quality solutions for the training set problems using the LLM: a popular recipe for reasoning tasks (Zelikman et al., 2022; Huang et al., 2023). Recent work has also attempted augmenting training set problems (Yue et al., 2024; Yu et al., 2024), but we limit our exploration to solution synthesis for existing problems in the benchmark.
Let’s solve this problem using Python code. <llm-code> discount_percent = 20 price_before_discount = 120 discount = discount_percent / 100 discount_amount = price_before_discount * discount price = price_before_discount - discount_amount price </llm-code> <llm-code-output> 96.0 </llm-code-output> So the new price of the lamp is 969696 dollars.
@@ 2.2.1 Default
We choose five representative examples of GSM8K and MATH to create the few-shot prompt for the respective datasets. For GSM8K, we use a mix of problems that require vanilla Python code and problems that are best solved using Python’s sympy library. For MATH, we compose a 5-shot prompt with examples from different subjects. To reflect this diversity of reasoning paths required for MATH, we choose a mix of problems that require code-based solutions, text-based solutions, and a combination of both. The prompts used for the two datasets are shown in Appendix B.6.
For GSM8K, we sample 128 solutions per training problem, which gets a training set coverage of 99.1%. For MATH, we sample 224 solutions per training problem, which only achieves a training set coverage of 80.1%. This difference in coverage reflects the difficulty of the MATH benchmark compared to GSM8K, which has been noted in previous work as well (Gou et al., 2024; Liao et al., 2024).
@@ 2.2.2 Subject-specific Prompting (Subj)
Could the diversity of mathematical topics in MATH be a reason for the low training set coverage with a single 5-shot prompt? To answer this question, we create subject-specific prompts for the seven subjects in the MATH benchmark, namely algebra, geometry, intermediate algebra, number theory, prealgebra, precalculus, and probability (See Table 10 in the appendix for the subject-wise split of MATH training data). The MATH benchmark also labels problems by their hardness level, with levels ranging from 1 to 5, where level 5 is the hardest. For creating subject-specific 5-shot prompts, we choose one example from each level for a given subject. For each of the seven prompts, we sample 32 solutions per problem and combine the data generated with all the prompts, which is equivalent to 32 x 7 = 224 solutions per problem. However, even with this fine-grained prompting, we only find a negligible gain in the training set coverage, though the total number of correct solutions increases by 14K (Table 2).
Combining this fine-tuning dataset with the earlier single default prompt dataset yields a training coverage of 85.1% for MATH, a boost of 5% absolute. But achieving this coverage required sampling almost 450 solutions per problem (224 + 224 = 448). Can we make the solution generation pipeline more efficient?
Question Lynne bought 7 books about cats and 2 books about the solar system. She also bought 3 magazines. Each book cost \$7 and each magazine cost \$4. How much did Lynne spend in all? Ground-Truth Text Solution Lynne bought a total of 7 + 2 = 9 books. The books cost Lynne 9 x 7 = \$63. For 3 magazines, Lynne spent 3 x 4 = \$12. In total, Lynne spent 63 + 12 = \$75 Masked Text Solution Lynne bought a total of 7 + 2 = M books. The books cost Lynne M x 7 = N. For 3 magazines, Lynne spent 3 x 4 = P. In total, Lynne spent N + P = Q
@@ 2.2.3 Masked Text Solution Prompting (Mask-Text)
GSM8K and MATH benchmarks come with ground-truth text solutions. Using these text solutions can, in theory, reduce the problem of code-interpreter solution generation to a translation problem from text to code. We initially experimented by prompting the LLM with:
@@
where $t_{i}$’s represent the text solution of representative problem $q_{i}$’s and $t^{\prime}$ represents the text solution of the problem $q^{\prime}$. Using the text solution in the prompt leads to a considerable increase in training set coverage. However, our manual analysis revealed that many solutions were shortcuts. E.g., trivial solutions such as print(ANSWER) or The answer is ANSWER where the ANSWER is copied from the text solution $t^{\prime}$ in the prompt. Our attempts to filter out these trivial solutions proved challenging as we ran into many creative ways in which the generated solution was cheating (see Figure 9 in Appendix).
@@ 2.3 Post-processing
While these post-processing steps can fix some of the syntactic errors, filtering semantically noisy, i.e., solutions that get to the right answer with flawed reasoning (Cobbe et al., 2021), is a much harder problem and beyond the scope of this work. Anecdotally, we find such solutions to be rare in our corpus. See Figure 11 in the Appendix for a sample semantically noisy solution.
(a) Naive Sampling
(b) Fair Sampling
Figure 4: Histogram of the number of solutions for problems in a 64K downsampled subset of MATH instances in OpenMathInstruct-1.
@@ 2.4 Data Selection
2.4.1 Fair vs Naive Downsampling
For a dataset like MATH, where problems can have very different difficulty levels, our solution generation strategy leads to a corpus where easier problems have a lot of solutions and harder problems have very few solutions (see Appendix A.3 for a detailed discussion on solution count). A naive strategy for downsampling treats every instance, i.e., problem-solution pair, as an equal. This problem-agnostic sampling perpetuates the imbalance of the original corpus, as seen in Figure 4(a). We propose a fair sampling alternate in which we iterate over all the problems round-robin and sample from unpicked solutions for each problem. This problem-dependent sampling ensures a more balanced representation of each problem in the downsampled dataset (see Figure 4(b)). Experimental results show that fair downsampling outperforms naive downsampling (Section 4.1.1).
@@ 3.1 Training Details
For all our experiments, including ablations, models of size 34B or smaller are trained for four epochs. A global batch size of 128 is used along with the AdamW optimizer with a weight decay of 1e-2 (Loshchilov and Hutter, 2019) and dropout (Hinton et al., 2012) of 0.1. We save one checkpoint per epoch for ablation experiments and two checkpoints per epoch for final model runs. The final checkpoint is created by averaging all the saved checkpoints. All experiments are performed using the NeMo toolkit [^9] https://github.com/NVIDIA/NeMo (Kuchaiev et al., 2019). For the full set of training hyperparameters, see Appendix B.1.
@@ 3.2 Evaluation Setup
We evaluate our models on the GSM8K and MATH benchmarks, which are also used to create the fine-tuning dataset. For ablation studies and hyperparameter selection, we create a validation set of 1K examples from the training set of GSM8K and MATH since both datasets lack an actual validation set. All the fine-tuned models are evaluated in the zero-shot setting. We use greedy decoding and self-consistency/majority voting (Wang et al., 2023) for evaluation. For majority voting, we found that using the lower temperature of 0.7 is beneficial compared to the data generation setup. We also deviate from the data generation setup by allowing the model to continue answering questions after code execution errors.

### v2

# spacing=nonfrench

@@ Abstract
Recent work has shown the immense potential of synthetically generated datasets for training large language models (LLMs), especially for acquiring targeted skills. Current large-scale math instruction tuning datasets such as MetaMathQA [1] and MAmmoTH [2] are constructed using outputs from closed-source LLMs with commercially restrictive licenses. A key reason limiting the use of open-source LLMs in these data generation pipelines has been the wide gap between the mathematical skills of the best closed-source LLMs, such as GPT-4, and the best open-source LLMs. Building on our proposed prompting novelty, the recent progress in open-source LLMs, and some brute-force scaling, we construct OpenMathInstruct-1, a high-quality math instruction tuning dataset with 1.8M problem-solution pairs. The dataset is constructed by synthesizing code-interpreter solutions for GSM8K and MATH, two popular math reasoning benchmarks, using the recently released and permissively licensed Mixtral model. Our best model, OpenMath-CodeLlama-70B, trained on a subset of OpenMathInstruct-1, achieves a score of 84.6% on GSM8K and 50.7% on MATH, which is competitive with the best gpt-distilled models. To support the open-source efforts, we have released our code, models, and the OpenMathInstruct-1 dataset under a commercially permissive license. [^1] Data and models are available at https://huggingface.co/collections/nvidia/openmath-65c5619de2ba059be0775014
@@ 1 Introduction
The huge development and inference costs associated with general-purpose large language models (LLMs) have led to the rise of smaller, task-specific LLMs. Recent work has proposed creating these domain/task-specific LLMs by generating high-quality synthetic data using powerful closed-source models such as GPT-3.5/4 [3] and training smaller models on the generated distillation data [4, 5, 6]. For mathematical reasoning, our task of interest, all the current state-of-the-art open-source models are gpt-distilled [7, 2, 8, 9]. However, model development recipes relying on proprietary models like GPT-4 can have serious limitations: (a) legal restraints on how the finetuned models can be used, [^2] https://openai.com/policies/terms-of-use (b) generating data with closed-source models is typically costlier than state-of-the-art open-source models, and (c) these recipes lack reproducibility as closed-source model behaviors can vary significantly over time [10].
@@
Figure 1: Training set coverage of Mixtral model generated solutions as a function of number of solutions sampled per problem (using temperature of 1.0 and top_p = 0.95). The statistics for the training set coverage of GPT-4 are from [8].
@@
For developing mathematical reasoning models, why are open-source models not used in place of closed-source models? To answer this, we compare GPT-4 with the Mixtral 8x7B model [11], one of the best open-source LLMs at mathematical reasoning in early 2024, by generating code-interpreter style solutions for two popular mathematical reasoning benchmarks, namely GSM8K [12] and MATH [13]. We use the metric training set coverage (TSC) to compare the models, where TSC measures the number of training problems for which any of the generated solutions leads to the ground truth answer (pass@k). Figure 1 shows the training set coverage (TSC) of the Mixtral model as a function of the number of sampled solutions. For the relatively easier GSM8K benchmark, the Mixtral model’s coverage catches up to GPT-4’s with almost 8x the number of solution samples. For the challenging MATH benchmark, even with 12x the number of solutions, the Mixtral model still has a lower TSC than GPT-4. This gap in the training set coverage reflects the distillation data quality and, hence, the quality of the final fine-tuned model. This explains the preference for GPT-4 in the current distillation pipelines for mathematical reasoning.
Bridging the coverage gap between GPT-4 and Open-source LLMs: We limit our investigation of open-source LLMs for synthesizing solutions to the Mixtral-base model due to (a) its strong performance on mathematical reasoning tasks compared to other open-source LLMs, and (b) its permissive license. [^3] https://mistral.ai/news/mixtral-of-experts/ As a first attempt, we use a brute-force approach of sampling several solutions per problem. However, this approach only scales logarithmically, limiting its effectiveness (Figure 1). Next, we explore the approach of targeted solution generation, where we write few-shot prompts focused on specific sections of the training data. Concretely, we write few-shot prompts for each mathematics subject in the MATH dataset and merge the synthesized solutions. The motivation is that these subject-specific few-shot prompts could better target the latent mathematical capabilities of these general-purpose LLMs. Unfortunately, we only find a marginal gain in TSC with this approach (Section 2.2.2). Finally, we utilize the fact that reference text solutions accompany mathematical benchmarks such as MATH and GSM8K. These reference solutions can aid the synthesis of code-interpreter style solutions. We show that using these reference solutions in our few-shot prompt with a slight modification substantially increases the coverage and, consequently, the performance of the fine-tuned model (Section 2.2.3 and 4.1.2).
Our solution synthesis experiments result in OpenMathInstruct-1, a collection of 1.8M problem-solution pairs. OpenMathInstruct-1 has a training set coverage of 93% for MATH and 99.9% for GSM8K. Table 1 shows that compared to previous mathematical reasoning fine-tuning datasets, OpenMathInstruct-1 is at least four times bigger, and, even more importantly, it is permissively licensed, allowing unrestricted usage by future work. To illustrate the quality of OpenMathInstruct-1, we train and release a range of models based on Mistral-7B [14], Llama 2 [15], and CodeLlama [16]. In particular, the CodeLlama-70B model fine-tuned on a subset of OpenMathInstruct-1, referred to as OpenMath-CodeLlama-70B, achieves a score of 84.6% on GSM8K and 50.7% on MATH. These scores are competitive with the current best GPT-distilled models. Finally, to support the open-source efforts in this direction, we have publicly released all our fine-tuned models, code, and the OpenMathInstruct-1 dataset, along with a further 6.6M incorrect sampled solutions under a commercially permissive license. [^4] The incorrect solution trajectories can be used to train verifier models [12, 17, 18].
@@
Dataset Size Generating LM (Permissive License)
Lila [19] 272K - ()
MathInstruct [2] 262K GPT-4 ()
MetaMathQA [1] 395K GPT-3.5 ()
MathCodeInstruct [7] 80K GPT-4 + Self ()
WizardMath* [20] 96K GPT-3.5 ()
ToRA* [8] 16K GPT-4 ()
@@ 2.1 Overview
Let $\mathcal{X}=\{(q_{1},a_{1}),\cdots,(q_{N},a_{N})\}$ be a typical mathematical reasoning training dataset, where $q_{i}$ and $a_{i}$ denote the $i^{\text{th}}$ question and answer respectively. Optionally, the training dataset may include reference text solution $t_{i}$, which illustrates a trajectory from $q_{i}$ to $a_{i}$ using mathematical principles. [^5] Both GSM8K and MATH have these text solutions. Besides the data, we assume access to a foundation LLM like Mixtral-base. The goal is to generate diverse, high-quality solutions for the training set problems using the LLM: a popular recipe for reasoning tasks [21, 22]. Recent work has also attempted augmenting training set problems [1, 2], but we limit our exploration to solution synthesis for existing problems in the benchmark.
@@ 2.2.1 Default
We choose five representative examples of GSM8K and MATH to create the few-shot prompt for the respective datasets. For GSM8K, we use a mix of problems that require vanilla Python code and problems that are best solved using Python’s sympy library. For MATH, we compose a 5-shot prompt with examples from different subjects. To reflect this diversity of reasoning paths required for MATH, we choose a mix of problems that require code-based solutions, text-based solutions, and a combination of both. The prompts used for the two datasets are presented in Appendix B.6.
For GSM8K, we sample 128 solutions per training problem, which gets a training set coverage of 99.1%. For MATH, we sample 224 solutions per training problem, which only achieves a training set coverage of 80.1%. This difference in coverage reflects the difficulty of the MATH benchmark compared to GSM8K, which has been noted in previous work as well [8, 9].
@@ 2.2.2 Subject-specific Prompting (Subj)
Could the diversity of mathematical topics in MATH be a reason for the low training set coverage with a single 5-shot prompt? To answer this question, we create subject-specific prompts for the seven subjects in the MATH benchmark, namely algebra, geometry, intermediate algebra, number theory, prealgebra, precalculus, and probability (See Table 11 in the appendix for the subject-wise split of MATH training data). The MATH benchmark also labels problems by their hardness level, with levels ranging from 1 to 5, where level 5 is the hardest. For creating subject-specific 5-shot prompts, we choose one example from each level for a given subject. For each of the seven prompts, we sample 32 solutions per problem and combine the data generated with all the prompts, which is equivalent to 32 x 7 = 224 solutions per problem. However, even with this fine-grained prompting, we only achieve a negligible gain in the training set coverage, though the total number of correct solutions increases by 14K (177K $\rightarrow$ 191K, see Table 2).
Combining this fine-tuning dataset with the earlier single default prompt dataset yields a training coverage of 85.1% for MATH, a boost of 5% absolute. However, achieving this coverage required sampling almost 450 solutions per problem (224 + 224 = 448). Can we make the solution generation pipeline more efficient?
Masked Text Solution Question
Lynne bought 7 books about cats and 2 books about the solar system. She also bought 3 magazines. Each book cost \$7 and each magazine cost \$4. How much did Lynne spend in all?
Ground-Truth Text Solution Lynne bought a total of 7 + 2 = 9 books. The books cost Lynne 9 x 7 = \$63. For 3 magazines, Lynne spent 3 x 4 = \$12. In total, Lynne spent 63 + 12 = \$75 Masked Text Solution Lynne bought a total of 7 + 2 = M books. The books cost Lynne M x 7 = N. For 3 magazines, Lynne spent 3 x 4 = P. In total, Lynne spent N + P = Q
@@ 2.2.3 Masked Text Solution Prompting (Mask-Text)
GSM8K and MATH benchmarks come with reference text solutions. Using these text solutions can, in theory, reduce the problem of code-interpreter solution generation to a translation problem from text to code. We initially experimented by prompting the LLM with:
@@
where $t_{i}$’s represent the text solution of representative problem $q_{i}$’s and $t^{\prime}$ represents the text solution of the problem $q^{\prime}$. Using the text solution in the prompt leads to a considerable increase in training set coverage. However, our manual analysis revealed that many solutions were shortcuts. E.g., trivial solutions such as print(ANSWER) or The answer is ANSWER where the ANSWER is copied from the text solution $t^{\prime}$ in the prompt. Our attempts to filter out these trivial solutions proved challenging as there are many creative ways in which the generated solutions were cheating (see Figure 12 in Appendix).
@@ 2.3 Post-processing
While these post-processing steps can fix some of the syntactic errors, filtering semantically noisy, i.e., solutions that get to the right answer with flawed reasoning [12], is a much harder problem and beyond the scope of this work. Anecdotally, we find such solutions to be rare in our corpus. See Figure 14 in the Appendix for a sample semantically noisy solution.
@@ 2.4 Data Selection
2.4.1 Fair vs. Naive Downsampling
For a dataset like MATH, where problems have divergent difficulty levels, our solution generation strategy leads to a corpus where easier problems have a lot of solutions and harder problems have very few solutions (see Appendix A.2 for a detailed discussion on solution count). A naive strategy for downsampling treats every instance, i.e., problem-solution pair, as an equal. This problem-agnostic sampling perpetuates the imbalance of the original corpus, as seen in Figure LABEL:fig:random_math. We propose a fair sampling alternate in which we iterate over all the problems round-robin and sample without replacement from the remaining solutions for each problem. This problem-dependent sampling ensures a more balanced representation of each problem in the downsampled dataset (see Figure LABEL:fig:fair_math). Experimental results show that fair downsampling outperforms naive downsampling (Section 4.1.1).
@@ 3.1 Training Details
For all our experiments, including ablations, models of size 34B or smaller are trained for four epochs. A global batch size of 128 is used along with the AdamW optimizer with a weight decay of 1e-2 [25] and dropout [26] of 0.1. We save one checkpoint per epoch for ablation experiments and two checkpoints per epoch for final model runs. The final checkpoint is created by averaging all the saved checkpoints. All experiments are performed using the NeMo toolkit [^9] https://github.com/NVIDIA/NeMo [27]. For the full set of training hyperparameters, see Appendix B.1.
@@ 3.2 Evaluation Setup
We evaluate our models on popular math reasoning benchmarks, namely GSM8K, MATH, GSM-Hard [24], SVAMP [28], TabMWP [29], ASDiv [30], and MAWPS [31]. For ablation experiments and hyperparameter selection, we create a validation set of 1K examples from the training set of GSM8K and MATH since both datasets lack an actual validation set. All the fine-tuned models are evaluated in the zero-shot setting. We use greedy decoding and self-consistency/majority voting [32] for evaluation. For majority voting, we found that using a lower temperature of 0.7 is beneficial compared to the data generation setup. We also deviate from the data generation setup by allowing the model to continue answer generation after code execution errors.